### v1

@@ 2.2 Explainable AI
Let the feature indices $\{1,\ldots,n\}$ be partitioned into two disjoint subsets: $A$ (variables of interest) and $C$ (complement). Also, let $\mathbf{x}=(\mathbf{x}_{A},\mathbf{x}_{C})$ with $\mathbf{x}_{A}\in\Omega_{A}$ and $\mathbf{x}_{C}\in\Omega_{C}$ and $P_{X_{C}}$ denote the marginal law of $X_{C}$. The partial dependence function of $\hat{f}$ on $\mathbf{x}_{A}$ is
@@
For continuous or ordinal variables, this variance quantifies the magnitude of influence of $x_{i}$ on the model output. However, for categorical variables, the notion of variance is not directly meaningful because categories lack a natural ordering. In such cases, a practical alternative sensitivity measure is the (normalized) range of the partial dependence values:
@@
Let $n$ be the number of features, and let $\{\mathbf{x}^{(j)}\}_{j=1}^{m}$ denote a dataset of $m$ instances. For each instance $j$, define a value-function
$v^{(j)}:2^{\{1,\dots,n\}}\;\to\;\mathbb{R},$
which assigns to every subset of features $V$ the expected model output when only the features in $V$ are observed for instance $j$:
$v^{(j)}(V)\;=\;\mathbb{E}\bigl[\hat{f}(\mathbf{X})\;\big|\;\mathbf{X}_{V}=\mathbf{x}^{(j)}_{V}\bigr$],
where the expectation is taken under a chosen background (reference) distribution for $\mathbf{X}$, choice that must be explicitly stated when reporting SHAP analyses [52].
The Shapley value of feature $i$ for instance $j$ is then defined by
$\Phi_{i}^{(j)}\;=\;\sum_{V\subseteq\{1,\dots,n\}\setminus\{i\}}\frac{\;|V|!\,(n-|V|-1)!}{n!}\;\bigl[v^{(j)}(V\cup\{i\})-v^{(j)}(V)\bigr],$
which corresponds to the average marginal contribution of feature $i$ across all coalitions $V$.
It follows that each model prediction admits the additive decomposition
$\hat{f}(\mathbf{x}^{(j)})\;=\;\phi_{0}^{(j)}\;+\;\sum_{i=1}^{n}\phi_{i}^{(j)},$
where $\phi_{i}^{(j)}=\Phi_{i}^{(j)}$ and $\phi_{0}^{(j)}=v^{(j)}(\varnothing)$. By construction, the SHAP values satisfy the core Shapley axioms (efficiency, symmetry, dummy, and additivity) and thus provide a model-agnostic, consistent, and fair feature attribution method [13].
@@ 4 Engineering Insights from Multidisciplinary Design Exploration
Each configuration is evaluated using a high-fidelity simulator to obtain target outputs, including fuel mass. A surrogate model is then trained on this dataset to emulate the costly simulator with substantially reduced computational cost. XAI techniques are applied to the surrogate to quantify the influence of individual design variables on fuel mass. These insights inform stakeholders during the co-design process, guiding refinements and highlighting trade-offs between competing objectives and constraints, such as energy performance and passenger capacity. This section demonstrates how our simulation workflow can be applied to the sequential design of a complex system while integrating explainability, extending our previous work [84] by handling mixed-variable input spaces, whereas most existing approaches are restricted to continuous variables.
@@ 4.1 Overall aircraft design and sizing of a Hybrid-Electric aircraft
This distributed electric propulsion concept, proposed by ONERA in 2019, was developed to support the European CleanSky2 initiative, which targets a 30% reduction in CO2 emissions by 2025. The DRAGON aircraft was designed to carry 150 passengers over a range of 2750 nautical miles at a cruise speed of Mach 0.78. While the configuration has previously been studied in the context of optimization for minimizing fuel mass [86, 88], there has been no in-depth investigation into the relationships between the design variables and fuel mass, an aspect that this study aims to address using explainability analysis [84].
@@
This problem features 10 continuous and 2 categorical input variables, in which the quantity of interest is fuel mass. The 10 continuous variables correspond to the geometrical and flight parameters of the aircraft as shown in Tab. 1. The first and second categorical variables correspond to the electric architecture with 17 levels and turboshaft layout level with 2 levels, respectively. The definition of the architecture variable is given in Tab. 2 and the definition of the turboshaft layout is given in Tab. 3.
@@ 4.2 Multidisciplinary design analysis
Designing complex engineering systems, such as aircraft, requires coordinating multiple interacting disciplines, including aerodynamics, structures, propulsion, and controls, where decisions in one area produce feedback that propagates through others and creates tightly coupled, often nonlinear behavior. To manage this complexity, eXtended Design Structure Matrix (XDSM) models offer a concise yet expressive way to represent modules, data flows, solver loops, and iteration order, making it easy to see which subsystems must be solved iteratively and which can be treated more loosely [60]. These abstract system models serve as blueprints that can be translated into executable Multidisciplinary Design Analysis (MDA) and optimization simulations in frameworks such as OpenMDAO, where each discipline is implemented as a component, drivers coordinate solvers and optimizers, and system-level derivatives or surrogates can be propagated for efficient gradient-based design exploration [47]. Together, XDSM and an MDAO improve traceability and reproducibility, guide solver and surrogate choices, and accelerate disciplined trade studies before committing to costly high-fidelity analyses.
For modeling electric architectures, it is more efficient to represent the architectural choices using two integer variables instead of one categorical variable. However, taking this approach expands the range of potential architectures beyond the initial 17 configurations. Yet, there are two important constraints to consider for these possible setups. The first constraint relates to the electrical connections between components: ensuring a certified electric architecture is crucial, and figuring out how to connect, for example, 8 motors to 6 generators is not straightforward. The second constraint is connected to the distributed propulsion system, especially the numerous propellers. Managing this system involves addressing a substantial number of potential failures in the electro-mechanical architecture, as for both stability and redundancy, not all electric connections are allowed. Consequently, to simplify the optimization problem and avoid introducing many constraints, the model uses a single categorical variable to represent the various feasible architectures.
@@ 4.4 Implementation of the simulator
In our workflow, the high-level XDSM representations are translated into executable Multidisciplinary Design Analysis (MDA) simulations, forming the computational core of the DRAGON co-design model. These simulations integrate aerodynamic, structural, propulsion, and electrical models within mission-analysis and sizing loops, ensuring that all estimated quantities reflect a physically consistent equilibrium achieved through multidisciplinary coupling and convergence towards consensus across disciplines.
The simulations rely on the open-source Future Aircraft Sizing Tool with Overall Aircraft Design (FAST-OAD) [27], enhanced with expert knowledge and recent extensions for hybrid-electric and distributed propulsion concepts. Built upon the OpenMDAO framework, FAST-OAD provides a modular structure where each discipline acts as a component, and where the data exchanges and solver couplings are explicitly defined. This setup enables flexible management of multidisciplinary interactions and clear control of convergence behavior. FAST-OAD is operated via a problem definition that enumerates design variables, constraints, solver drivers, and the set of discipline submodels to use. This modular configuration makes three operating modes straightforward: (i) single-point analyses for focused sampling, (ii) nested solver execution for fully coupled disciplinary convergence, and (iii) embedding FAST-OAD as the expensive blackbox evaluator inside surrogate-assisted global optimizers. Individual disciplines (geometry, aerodynamics, mass/weight, propulsion and power electronics, mission integration, electrical load-flow, etc.) can be swapped or augmented with custom submodels, enabling targeted fidelity management across the design space and straightforward incorporation of specific component behavior.
@@ 4.5 Surrogate modeling
A KPLS surrogate model is then trained over the 300-point dataset. This surrogate leads to a relatively small RMSE of 1.42% over the other 700-point dataset, making our surrogate appropriate for proper knowledge extraction. Data are available online [^1] https://github.com/ANR-MIMICO/EAAI_XAI.
@@ 5.5 Surrogate modeling
Note that our metrics are partial and subject to misinterpretation. For instance, the RMSE is an aggregated measure resuming average error and neither the maximal nor the median error, as such it gives little information on the error distributions, same for MCC which is less representative than the complete 4 values representing the confusion matrices or PVA which is based on a normal error assumption and on the idea that the predicted variance should be equal to the unknown error. We use RMSE to have an idea of the global regression precision of the model, MCC because it is a symmetric metric to quantify the confusion matrices and PVA because it consists on a simple metric that does not require any quantile computations. All results (RMSE, MCC and PVA) for the 16 surrogate models are reported in Tab. 5, where we have organized the models into the following categories: tree-based models (RF, GB, DT); kernel-based models (GP, RBF, SVM); linear and quadratic models (LR, QP); reduced basis models (RMTS, PCE); neighborhood-based models (kNN, IDW, CIEL); neural network models (MLP, GENN); and the transformer-based model (TabPFN). These various families are colored accordingly for more visibility, and the overall global aggregated results are in darker green when they have the best performances.
@@
Our results confirm that different surrogate models exhibit variable and complementary performances at the global scale, making them versatile and indicating promising future research towards a mixture of experts or ensemble approaches to combine various models. In particular, no method stands out at one of the top three for every metric, and no method falls into the bottom 4 for every metric.
@@
Ensemble methods have the disadvantage of requiring a large number of predictions to estimate uncertainties, thereby generating very discontinuous response surfaces, which can, in turn, limit their analytical capacity, interpretability, and gradient prediction, but they open up many avenues for research. However, and quite surprisingly, LR and its slightly more complex counterpart QP perform quite well on this problem even if they are simple polynomials of degree 1 and 2, respectively. Even if the latter behavior is hard to extrapolate to other problems, we argue that looking at and explaining the relationship of degree 1 or 2 is always good, as most global information is generally captured in such simple models, which can help detect relevant features really fast. Similarly, logic explanations such as abductive and contrastive explanations are known to be easy to compute based on a decision tree surrogate without any loss of generality [8]. Therefore, and even if RF and GP/RBF can outperform, respectively QP or DT we always recommend starting with these elemental, transparent surrogate models as they are the most efficient for a quick and first overview [49].
@@
To also study the local predictions, let’s examine more in-depth the qualitative behaviors of the three models that come with in-house uncertainty quantification. To do that, a spatial study comparing TabPFN, GP, and RF is provided in Fig. 12. In what follow, each variable has been fixed to its average value (or its integer part in the case of discrete variables) except for density, which varies between 0 and $1$, to illustrate the smooth aspect of the GP and the TabPFN model and the piecewise continuous aspect of the RF along with the poor uncertainty estimation associated with it, obtained by resampling over 100 estimators (default options), whereas the GP can be directly used for variance-based interpretability [53]. TabPFN variance is obtained with its quantile prediction and shows a clear non-smooth and non-interpretable behavior. In fact, TabPFN overall variance prediction has been shown to overestimate the uncertainty when computing the PVA, whereas the GP tends more to underestimates it, in particular here with repeated experiences and aleatoric intrinsic noise to model. This one-dimensional example illustrates the complementarity of predictive models and the importance of adapting the model to the studied scale at hand. Therefore, the figure shows that two models can agree on the expected mean prediction but still predict very different standard deviations, highlighting the complementarity of global and local analyses. The TabPFN model is really time-consuming in inference, but still achieved the lowest RMSE and second best in classification error, indicating superior ability to capture the nonlinear mapping from input parameters to sparsity, as it is also the case for RF and GP, which are well fitted for these low-data contexts. In fact, for this type of sparse and uniformly distributed data, it is generally preferable to use a GP or an RBF, as these models learn quickly and effectively capture nonlinear behaviors on small datasets [35]. Also, RF, as an ensemble approach, provides comparable results here, which can be explained by its ability to approximately reproduce the behaviors of nonlinear SABM. Gradient boosting and decision tree, coming slightly behind, also confirm the capacity of these methods to handle complex interactions among agents [95]. However, these results should be put in light of the approximation of such variance by conformal prediction, which is generally non-smooth. TabPFN’s variance approximation via interquantile range also performed reasonably, whereas both GP and RF tend to underestimate the error because of repeated experiences.
5.6 Explainable methods analyses and insights extraction

### v2

@@ 2.2 Explainable AI
Let the feature indices $N=\{1,\ldots,n\}$ be partitioned into two disjoint subsets: $A$ (variables of interest) and $C$ (complement). Also, let $\mathbf{x}=(\mathbf{x}_{A},\mathbf{x}_{C})$ with $\mathbf{x}_{A}\in\Omega_{A}$ and $\mathbf{x}_{C}\in\Omega_{C}$ and $P_{X_{C}}$ denote the marginal law of $X_{C}$. The partial dependence function of $\hat{f}$ on $\mathbf{x}_{A}$ is
@@
For continuous or ordinal variables, this variance quantifies the magnitude of influence of $x_{i}$ on the model output. However, for categorical variables, the notion of variance is not directly meaningful because categories lack a natural ordering. In such cases, a practical heuristic alternative sensitivity measure is the (normalized) range of the partial dependence values:
@@
The Shapley value, originally introduced in cooperative game theory [92], provides an axiomatic rule to fairly attribute a total payoff to individual players. Let $N$ denote the set of players (features) and $v:2^{N}\to\mathbb{R}$ a value function that assigns a payoff to each coalition $S\subseteq N$. The Shapley value of player $i\in N$ is
$\Phi_{i}(v)\;=\;\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\,(n-|S|-1)!}{n!}\,\bigl(v(S\cup\{i\})-v(S)\bigr),$ (5)
which equals the average marginal contribution of $i$ over all possible coalitions and satisfies the classical Shapley axioms (efficiency, symmetry, dummy, additivity).
The SHAP framework [67] adapts this principle to model explanations by treating features as players and the model prediction as the payoff. For a predictive model $\hat{f}$ and a specific instance $\mathbf{x}$, define the instance-specific value function
$v_{\mathbf{x}}(S)\;=\;\mathbb{E}\!\bigl[\hat{f}(\mathbf{X})\mid\mathbf{X}_{S}=\mathbf{x}_{S}\bigr],\qquad S\subseteq N,$
where the expectation is taken with respect to a chosen background distribution for $\mathbf{X}$ (the choice of which must be reported in practice). The SHAP value (feature attribution) of the feature $i$, for instance $\mathbf{x}$ is obtained by applying (5) to $v_{\mathbf{x}}$:
$\phi_{i}(\mathbf{x})\;=\;\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\,(n-|S|-1)!}{n!}\,\Bigl(v_{\mathbf{x}}(S\cup\{i\})-v_{\mathbf{x}}(S)\Bigr).$ (6)
By construction, the SHAP values provide the additive decomposition
$\hat{f}(\mathbf{x})\;=\;v_{\mathbf{x}}(\varnothing)\;+\;\sum_{i=1}^{n}\phi_{i}(\mathbf{x}),$
where $v_{\mathbf{x}}(\varnothing)$ is the prediction of the baseline (reference). To capture pairwise interactions, SHAP introduces interaction values $\phi_{i,j}(\mathbf{x})$ for $i\neq j$. One convenient symmetric definition is
$\phi_{i,j}(\mathbf{x})\;=\;\frac{1}{2}\sum_{S\subseteq N\setminus\{i,j\}}\frac{|S|!\,(n-|S|-2)!}{(n-1)!}\,\Bigl(v_{\mathbf{x}}(S\cup\{i,j\})-v_{\mathbf{x}}(S\cup\{i\})-v_{\mathbf{x}}(S\cup\{j\})+v_{\mathbf{x}}(S)\Bigr),$ (7)
which yields $\phi_{i,j}(\mathbf{x})=\phi_{j,i}(\mathbf{x})$ and decomposes each feature attribution as
$\phi_{i}(\mathbf{x})\;=\;\sum_{j=1}^{n}\phi_{i,j}(\mathbf{x}).$
The main (non-interaction) effect of feature $i$ is then commonly defined by
$\phi_{i,i}(\mathbf{x})\;=\;\phi_{i}(\mathbf{x})-\sum_{j\neq i}\phi_{i,j}(\mathbf{x}),$
so that the full decomposition across main effects and pairwise interactions is coherent and sums to the model prediction difference $\hat{f}(\mathbf{x})-v_{\mathbf{x}}(\varnothing)$.
Practical computation of the expectations $v_{\mathbf{x}}(S)$ requires a choice of background distribution and typically relies on approximations (sampling, model-specific fast algorithms, or conditional expectation estimators) due to the combinatorial number of subsets. These implementation choices affect both the numerical values and the interpretation of SHAP attributions and should therefore be reported alongside any SHAP-based analysis [67, 52, 29].
@@ 4 Engineering Insights from Multidisciplinary Design Exploration
Each configuration is evaluated using a high-fidelity simulator to obtain target outputs, including fuel mass. A surrogate model is then trained on this dataset to emulate the costly simulator with substantially reduced computational cost. XAI techniques are applied to the surrogate to quantify the influence of individual design variables on fuel mass. These insights inform stakeholders during the co-design process, guiding refinements and highlighting trade-offs between competing objectives and constraints, such as energy performance and passenger capacity. This section demonstrates how our simulation workflow can be applied to the sequential design of a complex system, integrating explainability, and extending our previous work [85] by handling mixed-variable input spaces, whereas most existing approaches are restricted to continuous variables.
@@ 4.1 Overall aircraft design and sizing of a Hybrid-Electric aircraft
This distributed electric propulsion concept, proposed by ONERA in 2019, was developed to support the European CleanSky2 initiative, which targets a 30% reduction in CO2 emissions by 2025. The DRAGON aircraft was designed to carry 150 passengers over a range of 2750 nautical miles at a cruise speed of Mach 0.78. While the configuration has previously been studied in the context of optimization for minimizing fuel mass [87], there has been no in-depth investigation into the relationships between the design variables and fuel mass, an aspect that this study aims to address using explainability analysis [85].
@@
This problem features 10 continuous and 2 categorical input variables, in which the quantity of interest is fuel mass. The 10 continuous variables correspond to the geometrical and flight parameters of the aircraft as shown in Tab. 1. The first and second categorical variables correspond to the electric architecture with 17 levels and turboshaft layout level with 2 levels, respectively. The architecture variable and the turboshaft layout are defined in Tables 2 and 3, respectively.
@@ 4.2 Multidisciplinary design analysis
Designing complex engineered systems, such as an aircraft, requires the coordinated work of several tightly coupled disciplines (aerodynamics, structures, propulsion, controls, etc.). A decision made in one subsystem, for example, a change in wing geometry, often triggers responses in others (drag, loads, engine demand), producing nonlinear interactions across multiple feedback loops. To tame this complexity, eXtended Design Structure Matrix (XDSM) diagrams provide a compact, yet expressive, graphical language for showing modules, data exchanges, solver loops, and execution order; these diagrams make it easy to spot which components must be solved iteratively and which can be treated more loosely [60]. Put simply, an XDSM is a schematic of the problem structure that helps teams see where coupling and iteration are required. These schematic descriptions can be converted into executable Multidisciplinary Design Analysis and Optimization (MDAO) workflows in frameworks such as OpenMDAO. In that setting, each discipline is wrapped as a component, drivers coordinate solvers and optimizers, and system-level derivatives or surrogate models are propagated to support efficient gradient-based exploration of the design space [47]. In plain terms, XDSM gives you the blueprint, and MDAO provides the assembly line to run experiments and optimizations automatically. Used together, they improve traceability and reproducibility, guide the choice of solvers and surrogate strategies, and enable faster, more disciplined trade studies before committing significant resources to costly high-fidelity simulations.
For modeling the electric architecture, it proves more effective to use two integer variables rather than a single categorical one to capture the design options. However, this choice significantly enlarges the space of possible architectures beyond the original 17 configurations. There are, nevertheless, two limitations that restrict this expanded domain. The first limitation is about how components are wired together because building a certifiable electrical topology is not simple. For example, deciding how to connect eight motors to six generators needs careful engineering. The second limitation concerns the distributed propulsion system and its many propellers. For safety and stability reasons, not every possible electrical connection is acceptable, and some arrangements would create too many potential failure modes. To keep the optimization manageable and avoid overcomplicating the constraint structure, the model ultimately reverts to a single categorical variable that encodes only the feasible and certified architectures.
@@ 4.4 Implementation of the simulator
In our workflow, the abstract XDSM diagrams are converted into executable multidisciplinary design analysis (MDA) simulations that constitute the computational backbone of the DRAGON co-design model. These simulations couple aerodynamic, structural, propulsion, and electrical submodels inside mission-analysis and sizing loops so that reported quantities correspond to a physically consistent operating point obtained by iterating across disciplines until mutual agreement is reached.
We implement these coupled analyses using the open-source Future Aircraft Sizing Tool with Overall Aircraft Design (FAST-OAD) [26], which we have extended with domain expertise and recent modules for hybrid-electric and distributed propulsion systems. FAST-OAD is built on the OpenMDAO infrastructure and follows a modular component paradigm: each discipline is encapsulated as a component, data transfers and solver couplings are made explicit, and the problem is defined by the set of design variables, constraints, driver(s), and chosen discipline submodels. This modularity makes the tool flexible in practice: it can run isolated single-point evaluations for sampling, execute nested solver sequences to obtain fully coupled convergence across subsystems, or serve as the computationally expensive black-box evaluator inside surrogate-assisted global optimization loops. Because individual discipline modules (geometry, aerodynamics, mass estimation, propulsion and power electronics, mission integration, electrical load-flow, etc.) can be replaced or enriched with user-provided submodels, the framework supports targeted fidelity control and straightforward incorporation of component-specific behavior across the design space.
@@ 4.5 Surrogate modeling
A KPLS surrogate model is then trained over the 300-point dataset. This surrogate leads to a relatively small RMSE of 1.42% over the other 700-point dataset, making our surrogate appropriate for proper knowledge extraction. Data are available online [^1] https://github.com/ANR-MIMICO/SMPT_XAI.
@@ 5.5 Surrogate modeling
Note that our metrics are partial and subject to misinterpretation. For instance, the RMSE is an aggregated measure resuming average error and neither the maximal nor the median error, as such it gives little information on the error distributions, same for MCC which is less representative than the complete 4 values representing the confusion matrices or PVA which is based on a normal error assumption and on the idea that the predicted variance should be equal to the unknown error. We use RMSE to have an idea of the global regression precision of the model, MCC because it is a symmetric metric to quantify the confusion matrices, and PVA because it consists of a simple metric that does not require any quantile computations. All results (RMSE, MCC and PVA) for the 16 surrogate models are reported in Tab. 5, where we have organized the models into the following categories: tree-based models (RF, GB, DT); kernel-based models (GP, RBF, SVM); linear and quadratic models (LR, QP); reduced basis models (RMTS, PCE); neighborhood-based models (kNN, IDW, CIEL); neural network models (MLP, GENN); and the transformer-based model (TabPFN). These various families are colored accordingly for more visibility, and the overall global aggregated results are in darker green when they have the best performances.
@@
Our results confirm that different surrogate models exhibit variable and complementary performances at the global scale, making them versatile and indicating promising future research towards a mixture of experts or ensemble approaches to combine various models. In particular, no method stands out as one of the top three for every metric, and no method falls into the bottom 4 for every metric.
@@
Ensemble methods have the disadvantage of requiring a large number of predictions to estimate uncertainties, thereby generating very discontinuous response surfaces, which can, in turn, limit their analytical capacity, interpretability, and gradient prediction, but they open up many avenues for research. However, and quite surprisingly, LR and its slightly more complex counterpart QP perform quite well on this problem, even if they are simple polynomials of degree 1 and 2, respectively. Even if the latter behavior is hard to extrapolate to other problems, we argue that looking at and explaining the relationship of degree 1 or 2 is always good, as most global information is generally captured in such simple models, which can help detect relevant features really fast. Similarly, logic explanations such as abductive and contrastive explanations are known to be easy to compute based on a decision tree surrogate without any loss of generality [8]. Therefore, and even if RF and GP/RBF can outperform, respectively QP or DT we always recommend starting with these elemental, transparent surrogate models as they are the most efficient for a quick and first overview [49].
@@
To also study the local predictions, let’s examine more in-depth the qualitative behaviors of the three models that come with in-house uncertainty quantification. To do that, a spatial study comparing TabPFN, GP, and RF is provided in Fig. 12. In what follow, each variable has been fixed to its average value (or its integer part in the case of discrete variables) except for density, which varies between 0 and $1$, to illustrate the smooth aspect of the GP and the TabPFN model and the piecewise continuous aspect of the RF along with the poor uncertainty estimation associated with it, obtained by resampling over 100 estimators (default options), whereas the GP can be directly used for variance-based interpretability [53]. TabPFN variance is obtained with its quantile prediction and shows a clear non-smooth and non-interpretable behavior. In fact, TabPFN overall variance prediction has been shown to overestimate the uncertainty when computing the PVA, whereas the GP tends more to underestimates it, in particular here with repeated experiences and aleatoric intrinsic noise to model. This one-dimensional example illustrates the complementarity of predictive models and the importance of adapting the model to the studied scale at hand. Therefore, the figure shows that two models can agree on the expected mean prediction but still predict very different standard deviations, highlighting the complementarity of global and local analyses. The TabPFN model is really time-consuming in inference, but still achieved the lowest RMSE and second best in classification error, indicating superior ability to capture the nonlinear mapping from input parameters to sparsity, as is also the case for RF and GP, which are well fitted for these low-data contexts. In fact, for this type of sparse and uniformly distributed data, it is generally preferable to use a GP or an RBF, as these models learn quickly and effectively capture nonlinear behaviors on small datasets [35]. Also, RF, as an ensemble approach, provides comparable results here, which can be explained by its ability to approximately reproduce the behaviors of nonlinear SABM. Gradient boosting and decision tree, coming slightly behind, also confirm the capacity of these methods to handle complex interactions among agents [95]. However, these results should be put in light of the approximation of such variance by conformal prediction, which is generally non-smooth. TabPFN’s variance approximation via interquantile range also performed reasonably, whereas both GP and RF tend to underestimate the error because of repeated experiences.
5.6 Explainable methods, analyses, and insights extraction